# A survey of degree-boundedness

### v1

@@ 5.3 String graphs and induced minors
Theorem 5.5 implies that the class of string graphs is degree-bounded. This implication can be obtained by using the Kővári-Sós-Turán Theorem [79] and the fact that every graph class with (strongly) sublinear separators actually has polynomial expansion, which was proven by Dvořák and Norin [48]. However, a much more direct proof is given by Fox and Pach in [56, Theorem 5]. As noted by Lee [88], the proof of Fox and Pach implies the following bound.
@@
In fact, it seems that the only missing piece is a generalization of [56, Lemma 1.7], which was originally proven in another paper by Fox and Pach [55].
@@ 5.4 Width parameters and a broader trend
As pointed out in Theorem 1.2, every graph class of bounded flip-width is degree-bounded due to a theorem of Toruńczyk [130]. This theorem implies that every class of bounded clique-width or bounded twin-width is degree-bounded, simply because such classes have bounded flip-width as well. However, the connections to degree-boundedness were proven even earlier for clique-width [69] and twin-width [15]. It is also not hard to show that many other classes of bounded width are degree-bounded, including classes of bounded tree-independence number [32], mim-width [76], and induced matching treewidth [89, 135].
These connections to width parameters actualy all suggest a more general trend. Let us say that a class of graphs $\mathcal{F}$ is weakly sparse if there exists an integer $t$ such that no graph in $\mathcal{F}$ has a subgraph isomorphic to $K_{t,t}$. This definition was introduced by Nešetřil, Ossona de Mendez, Rabinovich, and Siebertz [102] to help capture a common trend in structural graph theory. Namely, that properties of graph classes often come in pairs, with one property $P_{sparse}$ that is suitable in the sparse setting, and one property $P_{dense}$ which, informally, extends $P_{sparse}$ to the dense setting. (We think of a graph as being sparse if it has few edges and dense if it has many edges, informally.) These properties should agree on every weakly sparse graph class, that is, a class of graphs $\mathcal{F}$ should have property $P_{sparse}$ if and only if $\mathcal{F}$ has property $P_{dense}$ and $\mathcal{F}$ is weakly sparse.
The following theorem combines some well-known examples of this trend. The first three bullet points are equivalently stated as [102, Theorem 2.10], the fourth is from [61], the fifth is stated as [130, Theorem 6.3], and the sixth as [102, Corollary 2.3]. The final bullet point follows just from the definitions. The first three bullet points all correspond to width parameters which are closely related. While it may seem that there are an unbounded number of width parameters that one could consider, there is a real theory of width parameters which shows that (many, but not all) well-studied width parameters are instantiations of one abstract idea. This idea is discussed in [43], for instance. We also refer the reader to a paper of Gajarský, Pilipczuk, and Toruńczyk [62] for an explanation of how the following theorem is connected to transductions in the first-order logic of graphs.
Theorem 5.8.
@@
We note that the third bullet point originates with Gurski and Wanke [69], and that the third and fourth bullet points are proven using a theorem of Dvořák [45] as the main ingredient.
@@ Acknowledgments
We are grateful to Bryce Frederickson for pointing out the result of Golumbic and Goss [68, Theorem 4] in connection with degree-perfect graphs.

### v2

@@ 5.3 String graphs and induced minors
Theorem 5.5 implies that the class of string graphs is degree-bounded. This implication can be obtained by using the Kővári-Sós-Turán Theorem [85] and the fact that every graph class with (strongly) sublinear separators actually has polynomial expansion, which was proven by Dvořák and Norin [50]. However, a much more direct proof is given by Fox and Pach in [59, Theorem 5]. As noted by Lee [94], the proof of Fox and Pach implies the following bound.
@@
In fact, it seems that the only missing piece is a generalization of [59, Lemma 1.7], which was originally proven in another paper by Fox and Pach [58].
@@ 5.4 Width parameters and a broader trend
As pointed out in Theorem 1.2, every graph class of bounded flip-width is degree-bounded due to a theorem of Toruńczyk [140]. This theorem implies that every class of bounded clique-width or bounded twin-width is degree-bounded, simply because such classes have bounded flip-width as well. However, the connections to degree-boundedness were proven even earlier for clique-width [74] and twin-width [17]. It is also not hard to show that many other classes of bounded width are degree-bounded, including classes of bounded tree-independence number [34], mim-width [81, 141], and induced matching treewidth [1, 95, 146].
These connections to width parameters actually all suggest a more general trend. Let us say that a class of graphs $\mathcal{F}$ is weakly sparse if there exists an integer $t$ such that no graph in $\mathcal{F}$ has a subgraph isomorphic to $K_{t,t}$. This definition was introduced by Nešetřil, Ossona de Mendez, Rabinovich, and Siebertz [108] to help capture a common trend in structural graph theory. Namely, that properties of graph classes often come in pairs, with one property $P_{sparse}$ that is suitable in the sparse setting, and one property $P_{dense}$ which, informally, extends $P_{sparse}$ to the dense setting. (We think of a graph as being sparse if it has few edges and dense if it has many edges, informally.) These properties should agree on every weakly sparse graph class, that is, a class of graphs $\mathcal{F}$ should have property $P_{sparse}$ if and only if $\mathcal{F}$ has property $P_{dense}$ and $\mathcal{F}$ is weakly sparse.
The following theorem combines some well-known examples of this trend. The first three bullet points are equivalently stated as [108, Theorem 2.10], the fourth is from [17], the fifth is from [64], the sixth is stated as [140, Theorem 6.3], and the seventh is stated as [108, Corollary 2.3]. The final bullet point follows just from the definitions. We note that the third bullet point originates with Gurski and Wanke [74]. Shrubdepth was originally introduced as an analog of treedepth for dense graphs [67], although the first bullet point was proven later in [65, Lemma 2.12]. Finally, we also note that the fifth and sixth bullet points are proven using a theorem of Dvořák [47] as the main ingredient.
The first three bullet points all correspond to width parameters which are closely related to each other. While it may seem that there are an unbounded number of width parameters that one could consider, there is a real theory of width parameters which shows that (many, but not all) well-studied width parameters are instantiations of one abstract idea. This formalism is discussed in [45], for instance. We also refer the reader to a paper of Gajarský, Pilipczuk, and Toruńczyk [66] for an explanation of how the following theorem is connected to transductions in the first-order logic of graphs.
Theorem 5.8 ([17, 64, 65, 74, 108, 140]).
@@
  • $\mathcal{F}$ has bounded sparse twin-width if and only if $\mathcal{F}$ has bounded twin-width and is weakly sparse.
@@ Acknowledgments
We are grateful to Bryce Frederickson for pointing out the theorem of Golumbic and Goss [73, Theorem 4] in connection with degree-perfect graphs, to Tung Nguyen and António Girão for critical corrections to Sections 2.2 and 3.4, to Paul Seymour and Raphael Steiner for helpful improvements to Section 5.1, and to Shakhar Smorodinsky for discussions about the Zarankiewicz problem. We would also like to thank the referees for their careful comments that have improved the paper throughout.